OpenStudy (anonymous):

A Spartan missile and a Trojan missle are both flying at the same constant altitude. At time t, the spartan missile is at a point (16.1-7t+t^2 , 13t-2t^2} At time t, the trojan missile is at a point (26-13t+2t^2 , 23-5t+t^2). Do the missiles craash???

OpenStudy (dumbcow):

set coordinates equal to each other 2t^2 -13t+26 = t^2-7t+16.1 t^2 -6t + 9.9 = 0 solving for t using quadratic formula yields non-real solution so the 2 missiles never have equal x_coord at same time so they never crash

OpenStudy (anonymous):

Thank you makes a lot of sense!